OpenStudy (anonymous):

[WILL GIVE MEDALS]A ramp leads up to a building. The top of the ramp is 6 feet above the ground, and the bottom is 16 feet from the building as shown. What is the length of the ramp? Round your answer to the nearest tenth of a foot. A. 22.0 ft B. 16.6 ft C. 14.8 ft D. 17.1 ft

OpenStudy (anonymous):

I think its A.

OpenStudy (welshfella):

use the Pythagoras theorem if x is the length of the ramp x^2 = 16^2 + 6^2

OpenStudy (welshfella):

x^2 = 256 + 36 = 291 x = sqrt292

OpenStudy (anonymous):

So 17.1?

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

Oh ok thank you.

OpenStudy (welshfella):

yw remember the pythagoras theorem a^2 = b^2 + c^2 where ABC is a right angled triangle and a = the longest side ( called the hypotenuse)